The purchase price of a bond on March 3, 2019 is $400. If the yield rate is j2=11.1%j2=11.1%, how many days will it take for the dirty price to be at least $405.6. (1 point) The purchase price of a bond on March 3, 2019 is S400. If the yield rate is j2-11.1%, how many days will it take fo
Answer #1

Dirty Price = Price of Bond + accrued Interest

Accrued Interest = 405.6 - 400 = 5.6

Accrued Interest = Price * Rate * (Days/365)

Days = Accrued Interest*365/(Price*rate)

Days = 5.6*365/(400*11.1%) = 46 Days